Calculate the cell potential for the following reactions and predict whether the reaction will take place. If it will take place, write the cell notation.
1. Zn (s) + 2 Ag^+ (aq) → Zn^+2 (aq) + 2 Ag (s)
2. Cu (s) + Ni^+2 (aq) → Cu^+2 (aq) + Ni (s)
1
Expert's answer
2020-11-26T10:54:34-0500

The standard electrode potentials for Zinc, silver, copper and nickel ions are -0.763V, 0.800V, 0.337V and -0.230V respectively

1. Electrode potential = 0.763+ 0.800 = 1.563V

2. Electrode potential = 0.230+ 0.337 = 0.567V


Cell notation for 1. Zn/Zn2+ // 2Ag+ / 2Ag

Cell notation for 2. Cu/Cu2+ // Ni2+/Ni